OpenStudy (anonymous):

Michelle and her sister want to bake a cake. it takes michelle 3 hours and her sister 5 hours. how long will it take if they work together?

OpenStudy (anonymous):

1/3 + 1/5 = 1/x?

OpenStudy (anonymous):

that's all I know tbh XD I'm not sure what to do from here

hero (hero):

Working together formula \[\frac{M \times S}{M + S} = t\] Hours Michelle can bake a cake alone Hours her sister can bake a cake alone t = hours it take if they work together \[\frac{(3)(5)}{3 + 5} = t\] \[\frac{15}{8} = t\] \[\frac{8 + 7}{8} = t\] \[\frac{8}{8} + \frac{7}{8} = t\] \[1 + \frac{7}{8} = t\] So if they work together, it will take 1 and 7/8 hours @whalexnuker

OpenStudy (anonymous):

thanks but this is old :O i got it solved already!

hero (hero):

So approximately 1 hour and 53 minutes to bake a cake if both Michelle and her sister work together :)

hero (hero):

Yes, but I have given you a shortcut formula